9.3.11 Question Help In a random sample of 39 criminals convicted of a certain crime, it was determined that the mean length11/10/19 3:08 PM Math 1401 Section 087 chiniki fils Homework: Homework - Section 9.3 - Which Procedure Do I Use? Save Score:chiniki fils Math 1401 Section 087 11/10/19 3:09 PM Save Homework: Homework - Section 9.3 - Which Procedure Do T Use? 5 of 8
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 50

sample standard deviation = s = 11

sample size = n = 39

Degrees of freedom = df = n - 1 = 39 - 1 = 38

At 95% confidence level

\alpha = 1 - 99%

\alpha =1 - 0.95 =0.05

\alpha/2 = 0.025

t\alpha/2,df = t0.025, 38 = 2.024

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.024 * ( 11/ \sqrt 39)

Margin of error = E = 3.6

The 95% confidence interval estimate of the population mean is,

\bar x  ± E

= 50  ± 3.6

= ( 46.4, 53.6)

We can be 95% confident that the mean length of sentencing for the crime is between 46.4 and 53.6 months.